A circuit consists of a 100 Ω resistor, a 5 V battery, and an LED placed correctly and all in series. The voltage drop across the LED is 2 V. What is the current through the LED?
1
Expert's answer
2012-09-06T09:59:05-0400
Total voltage drop is 5 V. So, the voltage drop across the resistor is
Ur = 5
- 2 = 3 V.
The current in a circuit is I = Iled = Ir = Ur / R = 3 / 100 =
0.03 A.